OpenStudy (anonymous):

Find the surface area of a cone with a slant height of 11 and a radius of 15. Use 3.14 for π. Answer 1222 1220 1225 1227

OpenStudy (anonymous):



OpenStudy (anonymous):

3.14*15*11=LA

OpenStudy (anonymous):

B=pi*r^2

OpenStudy (anonymous):

I gave you what you need, now solve the problem. :)

OpenStudy (anonymous):

i did 3.14x15x11^2 and got 5,699.1 lol im confused

OpenStudy (anonymous):



OpenStudy (anonymous):

\[A= \pi r s + \pi r ^2\]

OpenStudy (anonymous):

do you know how to use the pythagoras theorem??

OpenStudy (anonymous):

is that the a^2 + b^2=c^2. If not then im not sure

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

oh sorry, you don't even need that to find the surface area...

OpenStudy (anonymous):

Its ok. I just dont get how to set the problem up to where i can get the right answer.

OpenStudy (anonymous):

you were given the radius and the slant height in your question....you were told to take the value of pi to be 3.14.... plug in the values into the formula to find the Surface area of the cone..

OpenStudy (anonymous):

I keep getting 7,771.5

OpenStudy (anonymous):

no...you are making a mistake.....i got a different answer.

OpenStudy (anonymous):

I know im making a mistake just dont no where. Can u give me the full equation so i can enter in my calcultor and see what i was doing wrong. Thanks

OpenStudy (anonymous):

\[A=\pi \times 15 \times 11 + \pi \times (15)^2\]

OpenStudy (anonymous):

ok i see thanks (:

OpenStudy (anonymous):

you are welcome....(:

OpenStudy (anonymous):

Oh one more if you have time. I need to Find the surface area of a square pyramid with a slant height of 21 and a base side measure of 18. I dont know the equation form for this one eathier. (:

OpenStudy (anonymous):

what do you get when you use the formula??

OpenStudy (anonymous):

Im not quite sure what the formula is?

OpenStudy (anonymous):

\[A=2bs + b^2\] b=base s=slant height

OpenStudy (anonymous):

\[A= 2(18)(21)+2(18)^2\] \[A=??\]

OpenStudy (anonymous):

@xr67cat